(10%) Problem 10: Suppose you connect a 7.25 μF and a 8.75 μF capacitor together 50% Part (a) What total capacitance will you make by connecting them in series in F? cosO cotanOasin acos( atanO acotan) sinh0 cosh0 tah cotanh0 Degrees Radians sinO tan Submit Hint I give up! Hints: 0% deduction per hint. Hints remaining: 3 Feedback: 0% deduction per feedback. 50% Part (b) What total capacitance will you make by connecting them in parallel in F?
Answer #1

Solution)

Given,

Capacitance, C1=7.25 uF

Capacitance, C2=8.75 uF

Part a)

Total Capacitance in series, Cs= (C1*C2)/(C1+ C2)

Substitute values,

Cs=(7.25*8.75)/(7.25+8.75)= 3.96 uF (Ans)

=======

Part b)

In parallel,

Cp= C1+C2=7.25+8.75= 16 uF (Ans)

=========